Write the following decimals as fractions. Reduce the fraction to the lowest form.
3.8

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

3.8 = 38 / 10
= 19 / 5

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
The given decimal is 3.8 
3.8 can be written as 38/10
 
Now, reducing it to simplest form = 38/10 = 19/5
 
Hope it helps!
Thank you